A. Determine the amount of the elements of SCI using the information provided.


DG Trading

Statement of Comprehensive Income

For the period ended, December 31, 20xx


-Advertising Expense - 5,000

-Bad Debts Expense - 10,000

-Commission Expense - 7,500

-Depreciaiton Expense - Furniture & Fixtures - 5,000

-Insurance Expense - 500

-Interest Expense - 2,000

-Light and Water Expense - 1,000

-Merchandise Inventory, December 31, 20xx - 30,000

-Merchandise Inventory, January 1, 20xx - 28,500

-Miscellaneous Selling Expense - 1,500

Office Salaries - 10,000

-Purchase Returns and Allowances - 1,500

-Purchases - 110,000

-Rent Expense - 7,500

-Sales - 237,500

-Sales Returns and Allowances - 2,500

-Sales Salaries - 15,000

-Taxes and Licenses - 2,500


B. Use the following to show your solution and underline twice your final answer.

1. Net Sales

2. Net Purchases

3. Cost of Sales

4. Gross Profit

5. Selling and Administrative Expenses

6. Net Income




1
Expert's answer
2021-09-09T11:36:35-0400

a) The Elements of an Income Statement

  • Revenue: Gross receipts earned by the company selling its goods or services.
  • Expenses: The costs to the company to earn the gross receipts.
  • Gains: Income from non-business-related transactions, such as selling a company asset.

Revenue = 237,500

Expenses = 5,000+ 10,000+ 7,500+5,000+500+2,000+1,000+1,500+10,000+7,500+2,500

Expenses=52,500

No gains

b) 1. Net sales is the sum of a company's gross sales minus its returns, allowances, and discounts.

Net Sales= Gross sales-Sales Returns-allowances-Discounts

Net sales =237,500-2,500-15,000

Net sales=220,000

2) Net purchases is defined as the gross amount of purchases made, less deductions for purchase discounts, returns, and allowances.

Net Purchases= Gross Purchases-Deductions for purchase discounts-returns-allowances

Net purchases=110,000-1500

Net purchases=108,500

3) Cost of Sales= beginning inventory + purchases - ending inventory

Cost of Sales=28,500+110000-30,000

Cost of Sales=108,500

4. Gross Profit = Revenue – Cost of Goods Sold

Gross Profit =220,000-108,500

Gross Profit =111,500

5. Selling and Administrative Expenses = 52,500

6. Net Income= Gross profit-selling and Administrative Expenses

Net Income=111,500-52,500

Net Income=59,000